For an input v(t)=5-2sin(100t)-cos(200t) ,the output of a full wave rectifier average is what?
Please give the explanation......

Grade:

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Hi Debanjan

v(t)=5-2sin(100t)-cos(200t)

for full wave rectifier

z(t)=mod (5-2sin(100t)-cos(200t))

v(t) is always greater than zero.

so z(t)=5-2sin(100t)-cos(200t)

avg. of sin and cosine function will be zero.

So the average of output is 5.

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..